# thanks: Corresponding Author

### v1

@@ IV Width of the molecule $\mathcal{M}_{\mathrm{c}}$
Here, we evaluate the width of the molecule $\mathcal{M}_{\mathrm{c}}$ by studying the decays $\mathcal{M}_{\mathrm{c}}\rightarrow\eta_{c}B_{c}^{+}$ and $\mathcal{M}_{\mathrm{c}}\rightarrow J/\psi B_{c}^{\ast+}$ which are dominant modes $\mathcal{M}_{\mathrm{c}}$. It is clear that both these modes are permitted channels for $\mathcal{M}_{\mathrm{c}}$. Indeed, the mass $\widetilde{m}=9712~\mathrm{MeV}$ of $\mathcal{M}_{\mathrm{c}}$ exceeds thresholds for these processes which amount to $9259~\mathrm{MeV}$ and $9435~\mathrm{MeV}$.
@@
In calculations the parameters $(M_{1}^{2},s_{0})$ in the channel of the molecule $\mathcal{M}_{\mathrm{c}}$ are chosen as in Eq. (14). The intervals for $(M_{2}^{2},s_{0}^{\prime})$ in the $B_{c}^{+}$ channel are chosen as
@@
i Channels $\widetilde{g}_{i}~(\mathrm{GeV}^{-1})\times 10^{2}$ $\Gamma_{i}~(\mathrm{MeV})$
@@ V Conclusions
Investigations carried out in the present work is a new step towards understanding of the internal structure and properties of the potential all heavy four-quark mesons. We have considered the scalar structures $bb\overline{b}\overline{c}$ and $cc\overline{c}\overline{b}$ organized as hadronic molecules $\mathcal{M}_{\mathrm{b}}=\eta_{b}B_{c}^{-}$ and $\mathcal{M}_{\mathrm{c}}=\eta_{c}B_{c}^{+}$. We have calculated their masses and evaluated decay widths by analyzed the dominant and some of subleading decay channels.
The masses and current couplings of these molecules have been calculated by means of QCD two-point sum rule method. Predictions $m=(15728\pm 90)~\mathrm{MeV}$ and $\widetilde{m}=(9712\pm 72)~\mathrm{MeV}$ obtained for the masses of $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{c}}$ have allowed us to determine their possible decay channels. In our studies we have distinguished the dominant and subleading decay mechanisms of these particles. The dominant mechanism is one in which all constituent quarks participate in producing of ordinary final-state mesons. For molecule $\mathcal{M}_{\mathrm{b}}$ breakdown to $\eta_{b}$ and $B_{c}^{-}$ mesons is the dominant process. The dominant channels of $\mathcal{M}_{\mathrm{c}}$ are the processes $\mathcal{M}_{\mathrm{c}}\rightarrow\eta_{c}B_{c}^{+}$ and $\mathcal{M}_{\mathrm{c}}\rightarrow J/\psi B_{c}^{\ast+}$. In the last decay $\mathcal{M}_{\mathrm{c}}$ falls to vector partners of the constituent mesons.
@@
All decay channels considered in this work have been explored using the three-point SR approach. This approach have permitted us to estimate the strong couplings $g_{i}$ and $\widetilde{g}_{i}$ at the vertices $\mathcal{M}_{\mathrm{b}}M_{1}M_{2}$ and $\mathcal{M}_{\mathrm{c}}M_{1}M_{2}$, where $M_{1}$ and $M_{2}$ are the final-state mesons. Our predictions $\Gamma\left[\mathcal{M}_{\mathrm{b}}\right]=(93\pm 17)~\mathrm{MeV}$ and $\Gamma\left[\mathcal{M}_{\mathrm{c}}\right]=(70\pm 10)~\mathrm{MeV}$ for the widths of the molecules $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{c}}$ mean that they may be interpreted as relatively broad structures. Note that numerous subleading processes form sizeable parts of these parameters.
@@
The hadronic molecules composed of four $b$ and $c$ quarks in various combinations were studied in Ref. Liu:2024pio . There, authors used the local gauge formalism to investigate the meson-meson interactions in such systems. In the scalar sector of this model, the molecular states rest above the relevant two-meson thresholds. Our findings for the scalar molecules $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{c}}$ are qualitatively consistent with this conclusion of Ref. Liu:2024pio . In this article the authors gave also information on parameter-dependent masses of axial-vector molecules $\Upsilon B_{c}^{-}$, $\eta_{b}B_{c}^{\ast-}$, and $\Upsilon B_{c}^{\ast-}$ which lie below the corresponding two-meson thresholds. It other words, these molecules can not dissociate to their ingredients, and in this sense, are stable structures. Of course, this does not mean that they are stable against the strong decays through annihilation mechanisms, which may lead to considerably broad structures even in these cases. Predictions of Ref. Liu:2024pio are interesting for understanding of the internal organizations and binding mechanisms of the fully heavy hadronic molecules, but need to be confirmed using alternative approaches including the sum rule method. This problem is beyond the scope of the present article, but eventually may be addressed in our future works.

### v2

@@ IV Width of the molecule $\mathcal{M}_{\mathrm{c}}$
Here, we evaluate the width of the molecule $\mathcal{M}_{\mathrm{c}}$ by studying its dominant modes $\mathcal{M}_{\mathrm{c}}\rightarrow\eta_{c}B_{c}^{+}$ and $\mathcal{M}_{\mathrm{c}}\rightarrow J/\psi B_{c}^{\ast+}$, as well as six channels generated by $c\overline{c}$ annihilation in $\mathcal{M}_{\mathrm{c}}$. It is clear that the both dominant decays are permitted channels for $\mathcal{M}_{\mathrm{c}}$. Indeed, the mass $\widetilde{m}=9712~\mathrm{MeV}$ of $\mathcal{M}_{\mathrm{c}}$ exceeds thresholds for these processes which amount to $9259~\mathrm{MeV}$ and $9435~\mathrm{MeV}$. Even in the lower limit $\widetilde{m}=9640~\mathrm{MeV}$ these decays are allowed channels of $\mathcal{M}_{\mathrm{c}}$. In this sense, it differs from the hadronic molecule $\mathcal{M}_{\mathrm{b}}$.
@@
In calculations the parameters $(M_{1}^{2},s_{0})$ in the channel of the molecule $\mathcal{M}_{\mathrm{c}}$ are chosen as in Eq. (16). The intervals for $(M_{2}^{2},s_{0}^{\prime})$ in the $B_{c}^{+}$ channel are
@@
i Channels $\widetilde{g}_{i}\times 10^{2}~(\mathrm{GeV}^{-1})$ $\Gamma_{i}~(\mathrm{MeV})$
@@ V Conclusions
Investigations carried out in the present work is a new step towards understanding of the internal structure and properties of the potential all heavy four-quark mesons. We have considered the scalar structures $bb\overline{b}\overline{c}$ and $cc\overline{c}\overline{b}$ organized as hadronic molecules $\mathcal{M}_{\mathrm{b}}=\eta_{b}B_{c}^{-}$ and $\mathcal{M}_{\mathrm{c}}=\eta_{c}B_{c}^{+}$. We have calculated their masses and evaluated decay widths by analyzing the dominant and subleading decay channels.
The masses and current couplings of these molecules have been calculated by means of QCD two-point sum rule method. The central values of the predictions $m=(15728\pm 90)~\mathrm{MeV}$ and $\widetilde{m}=(9712\pm 72)~\mathrm{MeV}$ mean that $\mathcal{M}_{\mathrm{b}}$ is close to $\eta_{b}B_{c}^{-}$ threshold, whereas $\mathcal{M}_{\mathrm{c}}$ locates a few hundred $\mathrm{MeV}$ above corresponding borders. In the lower limit for $m=15638~\mathrm{MeV}$ the structure $\mathcal{M}_{\mathrm{b}}$ can be interpreted as a bound state of the mesons $\eta_{b}$ and $B_{c}^{-}$. Contrary, $\mathcal{M}_{\mathrm{c}}$ does not form a bound state and is a broad resonance above relevant two-meson continuum.
Predictions for the masses of $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{c}}$ have allowed us to determine their possible decay channels. In our studies we have distinguished the dominant and subleading decay mechanisms of these particles. The dominant mechanism is one in which all constituent quarks participate in producing of ordinary final-state mesons. For molecule $\mathcal{M}_{\mathrm{b}}$ breakdown to $\eta_{b}$ and $B_{c}^{-}$ mesons is the dominant process. The dominant channels of $\mathcal{M}_{\mathrm{c}}$ are the processes $\mathcal{M}_{\mathrm{c}}\rightarrow\eta_{c}B_{c}^{+}$ and $\mathcal{M}_{\mathrm{c}}\rightarrow J/\psi B_{c}^{\ast+}$. In the last decay $\mathcal{M}_{\mathrm{c}}$ falls to vector partners of the constituent mesons.
@@
All decay channels considered in this work have been explored using the three-point SR approach. It has permitted us to estimate the strong couplings $g_{i}$ and $\widetilde{g}_{i}$ at the vertices $\mathcal{M}_{\mathrm{b}}M_{1}M_{2}$ and $\mathcal{M}_{\mathrm{c}}M_{1}M_{2}$, where $M_{1}$ and $M_{2}$ are the final-state mesons. Our predictions $\Gamma\left[\mathcal{M}_{\mathrm{b}}\right]=(93\pm 17)~\mathrm{MeV}$ and $\Gamma\left[\mathcal{M}_{\mathrm{c}}\right]=(70\pm 10)~\mathrm{MeV}$ for the widths of the molecules $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{c}}$ imply that they may be interpreted as relatively broad structures. Note that numerous subleading processes form sizeable parts of these parameters. In the scenario when $m=15638~\mathrm{MeV}$ the mesons $\eta_{b}$ and $B_{c}^{-}$ form the bound state $\mathcal{M}_{\mathrm{b}}$ which is, nevertheless, unstable hadronic molecule with $\Gamma\left[\mathcal{M}_{\mathrm{b}}\right]|_{\mathrm{l.m.}}=(49\pm 6)~\mathrm{MeV}.$
@@
The hadronic molecules composed of four $b$ and $c$ quarks in various combinations were studied in Ref. Liu:2024pio . There, authors used the local gauge formalism to investigate the meson-meson interactions in such systems. In the scalar sector of this model, the molecular states rest above the relevant two-meson thresholds. Our findings for the scalar molecules $\mathcal{M}_{\mathrm{b}}$ and $\mathcal{M}_{\mathrm{c}}$ are in part consistent with this conclusion of Ref. Liu:2024pio . In this article the authors gave also information on parameter-dependent masses of axial-vector molecules $\Upsilon B_{c}^{-}$, $\eta_{b}B_{c}^{\ast-}$, and $\Upsilon B_{c}^{\ast-}$ which lie below the corresponding two-meson thresholds. It other words, these molecules can not dissociate to their ingredients, and in this sense, are stable structures. Of course, this does not mean that they are stable against the strong decays through annihilation mechanisms, which may lead to considerably broad structures even in these cases. Predictions of Ref. Liu:2024pio are interesting for understanding the internal organizations and binding mechanisms of the fully heavy hadronic molecules, but need to be confirmed using alternative approaches including the sum rule method. This problem is beyond the scope of the present article, but eventually may be addressed in our future works.